OpenStudy (anonymous):

Solve: y divided by -6 + 5 = 9

OpenStudy (anonymous):

I dont know how to do this one at all

OpenStudy (aripotta):

does it look like this? 

OpenStudy (anonymous):

yeah the first one

OpenStudy (aripotta):

ok. what do you think the first step is? and honestly, i can say that this is much easier than you're thinking it is

OpenStudy (anonymous):

I dont know

OpenStudy (aripotta):

no ideas?

OpenStudy (aripotta):

remember 5c + 4 = −26? our first step was to subtract 4 from both sides...

OpenStudy (aripotta):

@jlongSwag27 c'mon, i know you know this :L

OpenStudy (anonymous):

Noo i dont know what to do

OpenStudy (aripotta):

le sigh

OpenStudy (aripotta):

subtract 5 from both sides.

OpenStudy (anonymous):

Okay so the answer is 24

OpenStudy (aripotta):

-24. see, you know what to do, you just want me to give you the answers

OpenStudy (anonymous):

haha no

OpenStudy (anonymous):

Solve: 2x + 10 + 26 = 212

OpenStudy (aripotta):

are you working on it?

OpenStudy (anonymous):

yeahhh

OpenStudy (anonymous):

@AriPotta Help me

OpenStudy (aripotta):

you know what to do though! what do you need help with? how far have you gotten?

OpenStudy (anonymous):

need help with this one Solve: 2x + 10 + 26 = 212 @AriPotta

OpenStudy (aripotta):

so what do you do first?

OpenStudy (anonymous):

Im not sure thats why im asking

OpenStudy (aripotta):

you have to combine like terms

OpenStudy (anonymous):

2x+36=212??

OpenStudy (aripotta):

yes, so now what

OpenStudy (anonymous):

answer is 88:)

OpenStudy (aripotta):

yep

OpenStudy (anonymous):

Solve: −3(x − 2) = 15 help me with this one plz.

OpenStudy (anonymous):

How do i start this one??

OpenStudy (aripotta):

you have to distribute the -3 into the parentheses

OpenStudy (anonymous):

So what would the answer be then??

OpenStudy (aripotta):

i'm not gonna just give you the answer lol

OpenStudy (aripotta):

what do you get after distributing?

OpenStudy (anonymous):

Plz haha, and im not sure for real though , i never did a 2 step equation like this

OpenStudy (aripotta):

−3(x − 2) = 15 -3x + 6 = 15 -3x = 9 x = -3

OpenStudy (anonymous):

thank you soo muchhh